Find the quadratic equation.
1. x = -1/3; x = 5/2

Answer by Edwin McCravy(20054)

It amounts to solving it backwards:

Start with the solutions:

x = -1/3;   x = 5/2

Clear of fractions:

3x = -1;  2x = 5

Get 0 on the right of each:

3x+1 = 0;  2x-5 = 0

Multiply equals by equals:

(3x+1)(2x-5) = 0

Foil that out:

6x²-15x+2x-5 = 0

6x²-13x+5 = 0

That's the quadratic equation.

Edwin